Please solve this question:
The ratio of boys to girls in Mr.Peter class is 3:4. On the last exam, 3/4 of the boys passed the exam and 4/5 of the girls passed. If a students exam paper is chosen at random, what is the probability that it has a passing grade?
also
How to do it on TI-84
Thanks

Answer by stanbon(75887)

The ratio of boys to girls in Mr.Peter class is 3:4.
# of boys = 3x
# of girls = 4x
Total # of boys and girls = 7x
----------------------
On the last exam, 3/4 of the boys passed the exam and 4/5 of the girls passed.
(3/4)(3x) = (9/4)x boys passed
(4/5)(4x) = (16/5)x girls passed
----------------------
If a students exam paper is chosen at random, what is the probability that it has a passing grade?
---
Total number who passed = [(9/4)+(16/5)]x = (45+56)/20 = (101/20)x
---
P(passed) = [101/20]x/7x = 101/140 = 0.7214
==================
Cheers,
Stan H.
==================